# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ 3.1 Background on Estimation
$\mu_{t}\in\arg\min_{g}\operatorname{E}_{P}[(Y-g(X))^{2}\mid T=t],t\in\{0,1\}$
Although many other advanced machine learning and causal inference methods have been developed based on advanced estimating equations [Nie and Wager, 2020, Kennedy, 2023, Oprescu et al., 2019, Wager and Athey, 2018, Semenova and Chernozhukov, 2021], or other machine-learning adaptations [Shalit et al., 2017, Shi et al., 2019], we will first instantiate our post-processing method with the $T$-learner and describe how it can be used with more advanced methods based on pseudo-outcome regression (such as the $DR$-learner [Kennedy, 2020, Semenova and Chernozhukov, 2021]).
Our meta-algorithm is based on post-processing CATE estimation (the $T$ or $DR-$ learner) with algorithms for multicalibration. Next, we describe multi-calibration and its prior use for universal adaptability in causal inference for the average treatment effect.
Universal Adaptability via Multicalibration
Recent work of [Kim et al., 2022] introduced the concept of universal adaptability in the context of causal inference. Much work on inference under (external) covariate shift assumes that the shift is known at the time of estimation. Instead, universal adaptability estimates the ATE from a dataset, with an estimate that incurs small bias on any downstream covariate distribution, within a broad class of unknown shifts. The work of [Kim et al., 2022] establishes the feasibility of universal adaptability via a connection to the notions of multicalibration/multiaccuracy, originally introduced in the literature on algorithmic fairness Hébert-Johnson et al. [2018]. Following this line of research, we show how multi-accuracy can be used, off-the-shelf, to address unknown (external and internal) shifts in the context of CATE.
@@
For $c(X)$ in a class of functions $\mathcal{C}$, a predictor $\tilde{p}:\mathcal{X}\rightarrow[0,1]$ is $(\mathcal{C},\alpha)$ multi-accurate if
@@
Specifically, we use the MCBoost algorithm [Pfisterer et al., 2021], pseudocode included in Algorithm 4. MCBoost [Pfisterer et al., 2021] takes as input a given initial predictor $p$, test-function class $\mathcal{C}$, approximation parameter $\alpha,$ and post-processing datasets $\mathcal{D}_{post}$ for calibration and validation. Later on in our meta-algorithms, to be concise we will refer to this as running $\textrm{MCBoost}(p,\mathcal{C},\alpha,\mathcal{D}_{post}).$ As a brief summary, MCBoost is a boosting procedure that proceeds via a series of auditing steps: given the initial predictor $p$, it then solves a least-squares problem over the calibration dataset to find a $c\in\mathcal{C}$ that maximizes correlation with the residuals. The predictor is then updated with a multiplicative-weights update based on the worst-case $c$. The process iterates until the total miscalibration or accuracy error drops below a stopping criterion. Next, we discuss the auditing step of this procedure in more detail.
Auditing
In practice, evaluation of the multi-calibration or multi-accuracy criterion over discrete subgroups is implemented via an auditing step that is reminiscent of twicing [Tukey et al., 1977]. (In the conditional moment literature, these audit test functions would be called instrument functions). That is, the algorithm often audits over real-valued functions $c(x)\colon\mathcal{X}\mapsto\mathbb{R}$. These can be connected to the subpopulation motivation by viewing $c(x)\colon\mathcal{X}\mapsto[0,1]$ as a relaxation of indicator functions; and real-valued functions as a rescaling of the former. (Later on, we will relate the real-valued weight functions directly to IPW weight functions (i.e. Riesz representers) in causal inference estimators). Given a predictor $p_{k}(x)$ at some iteration $k$ of the algorithm, the auditing step learns a test function $c$ that best correlates with the residual function $p_{k}(x)-y$. Auditing and postprocessing occurs in a different held-out dataset: we will refer to this as the post-processing dataset, including both calibration and validation sets. If the multi-accuracy criterion is not met for this test function $c\in\mathcal{C}$, the algorithm takes a boosting step and adds a multiplicative update with this test function. If the multi-accuracy criterion is met, the algorithm terminates.
@@
Let $\alpha>0,m\in\mathbb{N}$. Suppose $D_{\text{post}}\sim\mathcal{D}$ is a set of independent samples. A hypothesis $\tilde{p}:\mathcal{X}\rightarrow[0,1]$ passes $(\alpha)$-multiaccuracy auditing if for $h\in\arg\min\operatorname{E}_{{post}}[((\tilde{p}(x)-y)-h(x))^{2}],$
$|{{\operatorname{E}}}[(Y-\tilde{p}(X))h(x)]|\leq\alpha.$
@@
A reader in causal inference or econometrics may notice connections to conditional moment formulations. We expect that our later analysis, which is focused on multi-calibration/accuracy algorithms, also hold for adversarial formulations of conditional moments. (For example, Greenfeld and Shalit [2020] observes that adversarial moment conditions, in their case HSIC for independence of residuals, imply robustness to covariate shift).
(Meta)-Algorithm
We describe the meta-learner in Algorithm 1. We learn CATE estimates based on the $T-$learner (i.e. differencing outcome regression models). Then the multi-accurate CATE estimate is:
@@
$\operatorname{E}[\tilde{\tau}(X)].$
@@
$\hat{\mu}_{t}(x)\leftarrow\arg\min_{g\in\mathcal{G}}\operatorname{E}[(g-Y)^{2}\mid T=t],\text{ for }t\in\{0,1\}$
@@ 3.2 Warmup: Multicalibration, universal adaptability, and the ATE
As a precursor to introducing our method for robust CATE estimation with multi-calibration, we introduce properties of multi-calibration/multi-accuracy algorithms for estimation of the CATE and ATE. We begin with a result about how multi-calibrated predictors implies identification of the ATE under weaker functional specification conditions via regression adjustment. In the next subsection we show how these properties also can imply robust identification of the ATE and CATE under external covariate shifts.
Target-independent identification of the ATE under unconfoundedness
Our first exposition, when unconfoundedness holds, shows that multi-calibration/multi-accuracy can be viewed as finding a boosted predictor whose marginalization satisfies estimating equations for the average treatment effect. In addition, multi-calibration/multi-accuracy as an algorithmic scheme expands the functional complexity of the original predictor it is initialized with. Interestingly, regression adjustment with a multi-calibrated/multi-accurate predictor approximates the doubly-robust estimator for the ATE, and hence is consistent if either the original predictor is well-specified, the inverse propensity score is within the auditor function class, or if the prediction function is within the expanded function class output by multi-calibration/multi-accuracy.
@@
$\operatorname{E}[Y(1)-Y(0)]=\sum_{t\in\{0,1\}}\mathrm{E}\left[\frac{\mathbb{I}[T=t]}{e_{t}(X)}\left(Y-{\mu}_{t}(X)\right)+{\mu}_{t}(X)\right]$
@@
We can characterize multi-accurate learning with an auditor function class containing the inverse propensity score as an approximation of the AIPW estimate. As a note, the below statements hold up to an additional misspecification error (as shown in Kim et al. [2022]). Because the auditor function class is typically large (i.e. contains functions beyond the inverse propensity score), this is a “robust” way to conduct doubly-robust estimation.
Proposition 1 (Multi-accuracy implies robust estimation of the ATE).
Consider an auditor class $\mathcal{H}$ that is closed under affine transformation. Assume unconfoundedness holds. Consider the estimator $\operatorname{E}[\tilde{\tau}(X)]$ where $\tilde{\tau}(x)$ is the output of Algorithm 1 with auditor class $\mathcal{H}$, approximation parameter $\alpha$, initial outcome model estimators $\hat{\mu}_{1},\hat{\mu}_{0}$, and $\mathcal{D}_{post}$ from the same distribution as the data.
@@
Identification of the ATE under universal adaptability
We next recall robust identification properties of the ATE under potential violations of unconfoundedness. This was called “universal adaptability” in Kim et al. [2022] which studied missing data under unknown shifts, which directly implies robust identification for causal inference.
@@
$\operatorname{E}[Y(1)]$ $=\operatorname{E}[\operatorname{E}[Y\mathbb{I}[T=1]\mid X,U]\cdot\operatorname{E}[1/e^{*}(X,U)\mid X,U]]=\operatorname{E}[Y\mathbb{I}[T=1]\operatorname{E}[1/e_{1}(X,U)\mid X]]$
@@
(Note that $1/e(X)\neq\operatorname{E}[{1}/{e_{1}(X,U)}\mid X]$ due to Jensen’s inequality). Robust identification of the ATE follows from Equation 3, i.e. that $\mathcal{H}$ contains (approximately) $\operatorname{E}[{1}/{e_{1}(X,U)}\mid X].$ Essentially, assuming that the auditor function class contains the identifying (unknown weight), we can re-interpret the multi-accurate criterion as an approximation of adversarial IPW.
@@
Suppose that $W_{1}^{*}(X),W_{0}^{*}(X)\in\mathcal{H}.$ Run Algorithm 1 on $\mathcal{D}$ (possibly with unobserved confounders) over auditor function class $\mathcal{H}$ and outcome function class $\mathcal{G}$ to obtain $\tilde{\tau}(X)=\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)$. Then
@@
Of course, we have not gained identification for free: we cannot verify the assumption that $W_{1}^{*}(X),W_{0}^{*}(X)\in\mathcal{H}$ from observational data alone, just as we cannot test the unconfoundedness assumption from data alone. However, multi-calibration methods already work with quite flexible function classes, which could be nonparametric (RKHS, etc).
This is how multi-calibration confers general robustness to distribution shift, whether from the data generating process such as unobserved confounders, or from external covariate shifts at the time of deployment.

### v2

@@ 3 Background and related work
$\mu_{t}(x)\in\arg\min_{g(x)\in\mathcal{G}}\operatorname{E}_{P}[(Y-g(X))^{2}\mid T=t],t\in\{0,1\}$
Although many other advanced machine learning and causal inference methods have been developed based on advanced estimating equations [Nie and Wager, 2020, Kennedy, 2023, Oprescu et al., 2019, Wager and Athey, 2018, Semenova and Chernozhukov, 2021], or other machine-learning adaptations [Shalit et al., 2017, Shi et al., 2019], we will first instantiate our post-processing method with the $T$-learner and describe how it can be used with more advanced methods based on pseudo-outcome regression (such as the $DR$-learner; Kennedy 2020, Semenova and Chernozhukov 2021).
Our meta-algorithm is based on post-processing CATE estimation (the $T$- or $DR$-learner) with algorithms for multi-calibration. Next, we describe multi-calibration and its prior use for universal adaptability.
Universal Adaptability via Multicalibration.
Recent work of Kim et al. [2022] introduced the concept of universal adaptability. Much work on inference under (external) covariate shift assumes that the shift is known at the time of estimation. Instead, universal adaptability builds a prediction function from a source dataset whose average outcome estimate incurs small bias on any downstream covariate distribution, within a broad class of unknown shifts. The work of Kim et al. [2022] establishes the feasibility of universal adaptability via a connection to the notions of multi-calibration/multi-accuracy, originally introduced in the literature on algorithmic fairness [Hébert-Johnson et al., 2018]. Following this line of research, we show how multi-accuracy can be used, off-the-shelf, to address unknown (external and internal) shifts in the context of CATE.
@@
For $c(X)$ in a set of functions $\mathcal{C}$, a predictor $\tilde{p}:\mathcal{X}\rightarrow[-B,B]$ is $(\mathcal{C},\alpha)$ multi-accurate if
@@
Specifically, we use the MCBoost algorithm [Pfisterer et al., 2021], pseudocode included in Algorithm 4. MCBoost [Pfisterer et al., 2021] takes as input a given initial predictor $p$, test-function class $\mathcal{C}$, approximation parameter $\alpha,$ and post-processing datasets $\mathcal{D}_{post}$ for calibration and validation. For historical reasons, because the codebase was developed initially for multi-calibration which requires discretization of the interval $[0,1]$, we standardize the prediction range, under assumption 3. This pre-processing is completely without loss of generality. Later on in our meta-algorithms, to be concise we will refer to this as running $\textrm{MCBoost}(p,\mathcal{C},\alpha,\mathcal{D}_{post}).$ As a brief summary, MCBoost is a boosting procedure that proceeds via a series of auditing steps: given the initial predictor $p$, it then solves a least-squares problem over the calibration dataset to find a $c\in\mathcal{C}$ that maximizes correlation with the residuals. The process iterates until the total miscalibration or accuracy error drops below a stopping criterion. Next, we discuss the auditing step of this procedure in more detail.
Auditing.
The auditing step solves a subproblem of finding a subgroup with worst-case multi-accuracy error (bias).
@@
Let $\alpha>0$ and $\mathcal{H}$ be a test class of functions. Suppose $D_{\text{post}}\sim\mathcal{D}$ is a set of independent samples. A predictor $\tilde{p}:\mathcal{X}\rightarrow[-B,B]$ passes $(\alpha)$-multiaccuracy auditing if for $c^{*}\in\arg\min_{c\in\mathcal{C}}\operatorname{E}[((\tilde{p}(X)-Y)-h(X))^{2}],$
$|{{\operatorname{E}}}[(Y-\tilde{p}(X))c^{*}(X)]|\leq\alpha.$
In practice, evaluation of the multi-calibration or multi-accuracy criterion over discrete subgroups is implemented via regression, reminiscent of twicing [Tukey et al., 1977]. “Twicing” is a one-step boosting procedure: given a predictor function $p(x)$, fit the residual errors $c^{*}\in\arg\min_{c\in\mathcal{C}}\mathrm{E}[((Y-p(X))-c(X))^{2}]$ and return $p(x)+c^{*}(x)$ as the final prediction. (In the conditional moment literature, these audit test functions would be called instrument functions). That is, the algorithm often audits over real-valued functions $c(x)\colon\mathcal{X}\mapsto\mathbb{R}$. These can be connected to the subpopulation motivation by viewing $c(x)\colon\mathcal{X}\mapsto[0,1]$ as a relaxation of indicator functions; and real-valued functions as a rescaling of the former. (Later on, we will relate the real-valued weight functions directly to IPW weight functions (i.e. Riesz representers) in causal inference estimators).
The MCBoost algorithm incorporates these worst-case subgroup error functions to improve the predictor. Given a predictor $p_{k}(x)$ at some iteration $k$ of the algorithm, the auditing step learns a test function $c$ that best correlates with the residual function $p_{k}(x)-y$. For auditing over regression prediction functions, for step $k$, this is a regression problem: the auditing step solves $\min_{c\in\mathcal{C}}\mathrm{E}[((Y-p_{k}(X))-c(X))^{2}]$. The predictor is then updated with a multiplicative-weights update based on the worst-case test function $c^{*}(x)$, that is, the next predictor $p_{k+1}(x)$ is upweighted or downweighted based on the predicted error $c^{*}(x)$, $p_{k+1}(x)\propto e^{-\Delta_{c}\cdot c^{*}(x)/2}\cdot p_{k}(x),$ where $\Delta_{c}$ is the multi-accuracy error computed on a validation set, $\mathrm{E}_{V}[c^{*}(X)(Y-p_{k}(X))]$. Therefore the MCBoost procedure adjusts the original predictor based on the predicted errors ($c^{*}(x)$), adjusting more strongly in regions of higher error, as well as if the predictor incurs higher error on average. Auditing and postprocessing occurs in a different held-out dataset: we will refer to this as the post-processing dataset, including both calibration and validation sets. If the multi-accuracy criterion is not met for this test function $c\in\mathcal{C}$, the algorithm takes a boosting step and adds a multiplicative update with this test function. If the multi-accuracy criterion is met, the algorithm terminates.
@@
A reader in causal inference or econometrics may notice connections to conditional moment formulations. We expect that our later analysis, which is focused on multi-calibration/accuracy algorithms, also hold for adversarial formulations of conditional moments. (For example, Greenfeld and Shalit [2020] observes that adversarial moment conditions, in their case HSIC for independence of residuals, imply robustness to covariate shift). Recent advances in machine learning for conditional moment equations [Dikkala et al., 2020, Bennett and Kallus, 2023, Ghassami et al., 2022] typically develop min-max estimation algorithms that are unstable in practice; besides, the theory of conditional moment restrictions typically identifies finite-dimensional parameters rather than entire functions like the CATE. (See Section 6 for more extensive discussion of related work.)
Multi-accuracy vs. multi-calibration.
Throughout this paper, we have focused on a weaker multi-accuracy criterion [Kim et al., 2019], which is nonetheless sufficient for the robustness guarantees we seek for external validity and robustness to shifts. On the other hand, the stronger criterion of multi-calibration is somewhat more often studied in the literature. In this paper, we do not investigate the calibration properties of heterogeneous treatment effects, focusing instead on robustness to external shift. On the other hand, these are not completely unrelated: with slightly different notions, Wald et al. [2021] studies connections between multi-domain calibration and out-of-distribution generalization (without a focus on causal estimation). Other papers have investigated calibration for heterogeneous treatment effects [Xu and Yadlowsky, 2022], including orthogonalized isotonic regression [Van Der Laan et al., 2023].
4 Method
In this section, we develop and analyze our meta-algorithm that uses multi-accurate post-processing on the black-box regression models for a $T$-learner of the CATE. We first describe the meta-algorithm. To justify its use, we start with basic results focusing on ATE estimation. First we show how regression adjustment with multi-accurate outcome models approximates an augmented doubly-robust estimator (AIPW), in the absence of unobserved confounding. (This relates to estimation properties of our final approach). Second, we discuss the target-independent identification properties of ATE estimation, under the potential presence of unobserved confounding, which relates to our guarantees under unknown covariate shifts.
Identification implies we can write our target causal estimand in terms of functionals of the observational distribution alone. Estimating the regression-adjustment-identified functional with multi-accurate post-processing will result in certain equivalences with other estimation approaches, such as double robustness, with corresponding function classes. We establish results for the different settings we consider by combining these properties.
We then discuss the specific settings of unknown covariate shifts (external validity), observational/randomized data, and CATE estimation therein. Finally, a natural question remains: we established the properties of a multi-accurate $T$-learner, but what about more complex approaches? We provide extensions to recently studied pseudo-outcome regressions and prove approximate equivalence of a multi-accurate $T$-learner with a “weaker” multi-accurate $DR$-learner, suggesting that expanding the test function class can capture some of the improvements of a more complex estimation target.
4.1 (Meta)-Algorithm: Multi-accurate post-processed T-learner
We describe the meta-learner in Algorithm 1. We learn CATE estimates based on the $T$-learner (i.e. differencing outcome regression models). We post-process the outcome models with multi-accuracy. Then the multi-accurate CATE estimate is:
@@ 4 Method
$\mathrm{E}[\tilde{\tau}(X)]=\mathrm{E}[\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)].$ (3)
This represents estimating the ATE by imputing potential outcomes via regression: hence its name regression adjustment in the statistics literature, the “direct method” in the computer science literature, etc. Notably, the estimator does not explicitly use propensity scores. Nonetheless, we show how the robust test functions in multi-accurate auditing can nonetheless approximate inverse propensity score functions.
@@
$\hat{\mu}_{t}(x)\leftarrow\arg\min_{g(x)\in\mathcal{G}}\operatorname{E}[(g(X)-Y)^{2}\mid T=t],\text{ for }t\in\{0,1\}$
@@ 4.2 Warmup: Multicalibration, universal adaptability, and the ATE
We introduce properties of multi-calibration/multi-accuracy algorithms for estimation of the CATE and ATE. We begin by establishing that regression adjustment estimation with multi-accurate outcome predictors can approximate an AIPW estimator with certain nuisance functions. Later on, we combine these estimation properties with robust identification of the ATE and CATE under external covariate shifts.
Robust estimation of the ATE under unconfoundedness: regression adjustment with sufficiently post-processed multi-accurate outcome predictors = doubly-robust estimation.
As a warm-up, we first consider estimation when unconfoundedness holds. We show in the causal context, that multi-calibration/multi-accuracy can be viewed as finding a boosted predictor whose marginalization satisfies estimating equations for the average treatment effect. In addition, multi-calibration/multi-accuracy as an algorithmic scheme expands the functional complexity of the original predictor it is initialized with. Interestingly, regression adjustment with a multi-calibrated/multi-accurate predictor approximates the doubly-robust estimator for the ATE, and hence is consistent if either the original predictor is well-specified, the inverse propensity score is within the auditor function class, or if the prediction function is within the expanded function class output by multi-calibration/multi-accuracy.
@@
$\operatorname{E}[Y(1)-Y(0)]=\sum_{t\in\{0,1\}}\mathrm{E}\left[\frac{\mathbb{I}[T=t]}{e_{t}(X)}\left(Y-{\mu}_{t}(X)\right)+{\mu}_{t}(X)\right]$ (4)
@@
The approximation relies on conducting multi-accurate learning with an auditor function class containing the inverse propensity score. As a note, the below statements hold up to an additional misspecification error (as shown in Kim et al. [2022]). Because the auditor function class is typically large (i.e. contains functions beyond the inverse propensity score), this is a “robust” way to conduct doubly-robust estimation.
Proposition 1 (Multi-accuracy implies robust estimation of the ATE via regression adjustment).
Suppose Assumptions 1, 2 and 3. Consider an auditor class $\mathcal{H}$ that is closed under affine transformation. Assume unconfoundedness holds. Consider the estimator $\operatorname{E}[\tilde{\tau}(X)]$ where $\tilde{\tau}(x)$ is the output of Algorithm 1 with auditor class $\mathcal{H}$, approximation parameter $\alpha$, initial outcome model estimators $\hat{\mu}_{1}(x),\hat{\mu}_{0}(x)$, and $\mathcal{D}_{post}$ from the same distribution as the data.
@@
Although identification of the ATE with doubly-robust estimator is standard, note that our estimator that uses multi-accurate outcome models does not explicitly estimate the propensity scores. Nonetheless, the novelty of Prop. 1 is that averaging multi-accurate outcome models, because multi-accuracy implies $\mathrm{E}[(Y-\tilde{\mu}_{t}(X))h(X)]\leq\epsilon,\forall h\in\mathcal{H}$ leads to direct approximation of a doubly-robust estimator with the multi-accurate regression-based estimator $\mathrm{E}\left[\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)\right]$, under conditions on how the test function class $\mathcal{H}$ relates to inverse propensity score functions. Because this is a novel perspective, we establish it for ATE estimation first before showing how it applies to the robust covariate shifts that motivate our method.
@@
Although the beneficial estimation properties of AIPW are well-established [Robins et al., 1994, Kennedy, 2016], recall that our multi-accurate T-learner was not explicitly constructed as an AIPW estimator. It is therefore an example of an outcome-regression based approach that nonetheless satisfies doubly-robust estimating equation properties, in the spirit of [Bang and Robins, 2005], but algorithmically different. This additional perspective on how multi-accurate learning may robustly approximate a doubly-robust estimator is new in our paper, beyond previous analysis of universal adaptability. It also relates in spirit to recent work studying connections between regression adjustment and IPW methods, as [Chattopadhyay and Zubizarreta, 2023] does for linear regression. We build on this characterization in Proposition 3 to relate our later developments to doubly-robust CATE estimators.
Robust target-independent identification and estimation of the ATE under universal adaptability.
Next we show how estimation with multi-accurate learning can robustly identify and estimate the ATE under potential violations of unconfoundedness. This is a re-interpretation of “universal adaptability” in Kim et al. [2022] which studied missing data under unknown shifts, which directly implies robust identification for causal inference.
@@
$\operatorname{E}[Y(1)]$ $=\operatorname{E}[\operatorname{E}[Y\mathbb{I}[T=1]\mid X,U]\cdot\operatorname{E}[1/e^{*}_{1}(X,U)\mid X,U]]=\operatorname{E}[Y\mathbb{I}[T=1]\operatorname{E}[1/e_{1}(X,U)\mid X]]$
@@
(Note that $1/e_{1}\neq\operatorname{E}[{1}/{e_{1}(X,U)}\mid X]$ due to Jensen’s inequality). Robust identification of the ATE follows from Equation 5, i.e. that $\mathcal{H}$ contains (approximately) $\operatorname{E}[{1}/{e_{1}(X,U)}\mid X].$ Essentially, assuming that the auditor function class contains the (unknown) identifying weight, we can re-interpret the multi-accurate criterion as an approximation of adversarial IPW. We summarize this in the following corollary.
@@
Suppose Assumptions 3 and 2. Suppose that $W_{1}^{*}(X),W_{0}^{*}(X)\in\mathcal{H}.$ Run Algorithm 1 up to $(\alpha)$-multiaccuracy on $\mathcal{D}$ (possibly with unobserved confounders) over auditor function class $\mathcal{H}$ and outcome function class $\mathcal{G}$ to obtain $\tilde{\tau}(X)=\tilde{\mu}_{1}(X)-\tilde{\mu}_{0}(X)$. Then
@@
Of course, we have not gained identification for free: we cannot verify the assumption that $W_{1}^{*}(x),W_{0}^{*}(x)\in\mathcal{H}$ from observational data alone, just as we cannot test the unconfoundedness assumption from data alone. However, multi-accuracy/multi-calibration methods already work with quite flexible function classes, which could be nonparametric (RKHS, etc).
This is how multi-accuracy confers general robustness to distribution shift, whether from the data generating process such as unobserved confounders, or from external covariate shifts at the time of deployment.